If 0.81 N is force of repulsion between two identical charges separated by a distance 0.1 m apart. The value of each charge in C will be


1
Expert's answer
2021-04-26T17:12:04-0400

Electrostatic force is F=0.81 N

Distance between two charges is d=0.10 m

Coulomb's constant is "k=9\times 10^9 N.m^2/C^2"

The electrostatic force between two charges is

"F=\frac{kQ^2}{d^2}"

The each charge is

"Q=\sqrt{\frac{Fd^2}{k}}=\sqrt{\frac{(0.81)(0.10)^2}{9\times 10^9}}=9.50\times 10^{-7} C"